# Resonant Photon-Axion Mixing Driven by Dark Matter Oscillations

### v1

@@ Abstract
A distinctive consequence of this driven mixing is the generation of polarization asymmetries between propagation modes. In particular, the stochastic crossing of Floquet resonances induces circular polarization with a characteristic variance set by the resonance strength. Since the axion-induced modulation affects the two circular polarization states with opposite signs, Floquet resonances generically break the symmetry between them. As a result, photon propagation through an axion DM background becomes intrinsically polarization-sensitive, even for initially unpolarized states.
In particular, repeated or stochastic crossings of Floquet resonances lead to a net circular polarization with vanishing mean but nonzero variance. This stochastic circular polarization arises from coherent, unitary mode mixing and is therefore qualitatively distinct from polarization generated by absorption or scattering processes. The magnitude of the variance is directly controlled by the resonance strength, thereby providing a direct probe of the driven mixing mechanism. Importantly, this polarization signature does not rely on fine-tuned source properties or detailed environmental modeling. It follows directly from the existence of Floquet resonances in photon propagation through a coherent axion background.
Constraints from the balazar 3C 279.– We validate this theoretical framework with full numerical solutions and apply it to the optical polarimetry of the blazar 3C 279 as an example. The parametric resonance breaks the symmetry between circular polarization states, inducing a stochastic signal in the blazar jet. Detailed configurations are provided in the supplementary material. To derive constraints on the axion–photon coupling $g_{a\gamma}$, we adopt a conservative statistical criterion: the standard deviation of the axion-induced circular polarization $\sigma_{\mathrm{std}}$ must not exceed the characteristic observational uncertainty of the blazar 3C 279 ($\sigma_{\Pi_{C}}\approx 0.30\%$ [Liodakis:2021els]). This ensures consistency between the predicted stochastic signal and the measured noise level. For a given astrophysical source, the axion phase $\varphi_{a}$ is effectively constant over observational timescales. However, since its value is a priori unknown, the axion-induced circular polarization behaves observationally as a stochastic contribution with zero mean. Even when the axion-induced circular polarization exhibits a bimodal distribution for certain masses, its sign and magnitude remain unpredictable due to the unknown axion phase, and the variance provides a conservative measure of its observational impact. Requiring its variance not to exceed the instrumental sensitivity therefore provides a conservative consistency bound. Fig. 2 shows the illustrative constraint on parameter space, with CAST bounds [CAST:2017uph] included for comparison.
For $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, we obtain conservative limits as $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$. The distinctive band-like structure of the exclusion contour is a signature of the Floquet parametric resonance mechanism, reflecting the discrete nature of the resonance condition. Two features emerge: (1) The conversion efficiency, affected by the Bessel functions $J_{n}(\Delta_{\text{DM}}/m_{a})$, is maximized when the effective momentum mismatch is bridged by integer harmonics of $m_{a}$. The sharp dips in the exclusion limit correspond to ”sweet spots” where this resonance condition is optimally satisfied within the jet profile. (2) As $m_{a}$ varies, the dominant contribution shifts between different resonance orders $n$. Weaker constraints between bands reflect off-resonance regimes, where the system approaches the adiabatic limit, resulting in reduced circular polarization production.
The constraints in Fig. 2 show that axion-induced parametric resonance leads to verifiable physical consequences rather than negligible perturbative effects. Although the current limit region is set by the specific noise floor of 3C 279, the concrete limit validates the physical relevance of our framework.
@@
Figure 2: Exclusion contours on the axion–photon coupling $g_{a\gamma}$ derived from 3C 279 optical polarimetry. The shaded region indicates parameters excluded by the non-observation of excess circular polarization. The resulting exclusion exhibits a distinctive band structure that directly reflects the discrete Floquet resonance condition $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, providing a characteristic signature of the driven mixing mechanism.
@@
Conclusions.– Wave propagation in a coherently oscillating background is intrinsically a periodically driven problem. We have shown that photon propagation in axion DM therefore exhibits Floquet dynamics, leading to resonant photon–axion mode conversion beyond the conventional static or adiabatic mixing paradigm. This driven resonance arises when the dispersion mismatch between photons and axions is compensated by integer harmonics of the axion oscillation frequency, $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, enabling coherent conversion even far from the standard level-crossing condition.
@@
The resulting Floquet resonance represents a distinct class of propagation phenomena. Unlike traditional resonant conversion, which is localized to specific spatial regions where plasma and axion masses coincide, the driven resonance identified here originates from temporal phase matching induced by the oscillating DM background. As a consequence, resonant mixing can persist over extended propagation distances, allowing small conversion amplitudes to accumulate coherently and imprint robust signatures on photon polarization.
The significance of this mechanism depends on the stability of the driven system against environmental fluctuations. In rapidly varying media, strong gradients can disrupt the phase coherence required for Floquet mixing, suppressing the resonance. In sufficiently smooth environments, however, the driven propagation effect can dominate photon–axion conversion, opening a coherence-based window for probing axion dark matter that is complementary to localized, high-density resonance searches.
More broadly, the Floquet framework developed here is not restricted to a single axion species. Because the resonance condition relies on the periodic structure of the background rather than its microscopic identity, analogous driven conversion phenomena may arise in multi-axion systems or more complex dark-sector environments. Our results therefore establish a general foundation for studying resonant wave propagation in coherently oscillating media, revealing a previously unexplored regime of dark-matter-induced photon conversion.
@@ Acknowledgements.
Acknowledgments.– The authors would like to thank the computing cluster of School of Fundamental Physics and Mathematical Sciences at Hangzhou Institute for Advanced Study, UCAS for providing the computational resources used in this work. This work is supported by the National Key Research and Development Program of China Grant No. 2020YFC2201502, and the National Natural Science Foundation of China under Grants No. 12475065, 12250010, 12447131, 12447105, and 12575113.
@@
This Supplementary Material (SM) contains additional calculation and derivation in support of the results presented in this work. First, we give a full derivation of the mixing matrix $\mathcal{M}$. Then, we present additional details regarding the derivation of the two-level systems. Next, we discuss the relation between our work and parametric instability. Finally, we present a numerical analysis and apply the formalism to the relativistic jet of blazar 3C 279.
@@
For a weakly magnetized cold plasma, the dielectric tensor is isotropic, $\bm{\epsilon}=(1-\frac{\omega_{\mathrm{pl}}^{2}}{\omega^{2}})\mathbb{I}$, where $\omega_{\mathrm{pl}}=\sqrt{4\pi\alpha n_{e}/m_{e}}$ is the plasma frequency and $\omega$ is the energy of the photon–axion system. In the radiation gauge, longitudinal plasma oscillations decouple from the propagating electromagnetic modes. In Fourier space, the induced plasma current therefore contributes an isotropic dispersion term $\mathbf{J}_{\mathrm{ind}}(\omega)=-\omega^{2}(\mathbb{I}-\bm{\epsilon})\mathbf{A}(\omega)=-\omega_{\mathrm{pl}}^{2}\mathbf{A}(\omega)$ to both photon polarizations, which can be absorbed into an effective photon mass.
@@
$\mathcal{M}(z)=\begin{pmatrix}\Delta_{\perp}&\Delta_{F}&0\\ \Delta_{F}^{*}&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$, (S13)
where $\Delta_{\perp}=\Delta_{\parallel}\equiv-\omega_{\mathrm{pl}}^{2}/(2\omega)$, $\Delta_{a}\equiv-m_{a}^{2}/(2\omega)$, and $\Delta_{\mathrm{B}}\equiv g_{a\gamma}B/2$. The $\Delta_{F}$ term originates from the axion-gradient-induced birefringence and is formally analogous to the Faraday rotation term. Explicitly, it is given by
@@
In the mixing matrix, we have neglected the absorption effects of photons, the contribution of the cosmic microwave background’s energy density to the dispersion relation, and the QED vacuum birefringence effect.
@@ II Effective Two-Level systems
where $\Delta_{\pm}=\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi$. Neglecting the QED birefringence effect, the evolution for the photon state can be formally written as
@@
Repeating the Fourier harmonic expansion as above, one would also obtain the same resonance condition and Rabi oscillations. Although axion-induced frequency modulation has been noted in earlier studies [McDonald:2019wou], it is often incorporated as a slowly varying phase or an effective modification of the refractive index. Such treatments effectively average over the harmonic structure of the modulation and therefore do not capture the Floquet sidebands responsible for resonant mode mixing.
@@ III.2 Regime I: Parametric Instability (The Mathieu Limit)
where $\epsilon\sim g_{a\gamma}a_{0}m_{a}/\omega$. The defining feature of this regime is that the periodic axion background modulates the effective frequency squared of the oscillator. This modulation couples the positive-frequency ($e^{-i\omega t}$) and negative-frequency ($e^{+i\omega t}$) modes. When $m_{a}\simeq 2\omega$, the system enters an instability band where the field amplitude grows exponentially. This corresponds to the stimulated decay of the axion field into photon pairs, a genuine parametric instability where the background acts as an energy pump.
@@ IV Numerical analysis
To numerically verify the analytical framework, we solve the Eq. (S16) incorporating the mixing matrix $\mathcal{M}$ for a concrete astrophysical environment. We model the relativistic jet of blazar 3C 279, focusing on the unpolarized optical R-band photons ($E_{\gamma}\approx 1.9\,\mathrm{eV}$) with a representative parameter set: an axion mass $m_{a}=10^{-22}\,\mathrm{eV}$ and a coupling constant $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$. The radiation zone, located $r_{0}=0.5\,\mathrm{pc}$ from the central engine, is characterized by a magnetic field $B_{0}=0.5\,\mathrm{G}$ and electron density $n_{e,0}=5\times 10^{3}\,\mathrm{cm}^{-3}$ [A:2025wjt]. The jet is modeled as a conical outflow with a Doppler factor $\delta\simeq 13$ [A:2025wjt], assuming scaling profiles $B(r)\propto r^{-1}$ and $n_{e}(r)\propto r^{-2}$ over a total length $L_{\rm jet}=1\,\mathrm{kpc}$. To capture the fine-grained evolution of the mixing process, the calculation is performed over a domain size ranging from $5\times 10^{-7}\,\mathrm{pc}$ to $10^{-3}\,\mathrm{pc}$. To isolate and illustrate the resonance structures dictated by the plasma profile, we adopt a uniform DM density $\rho_{\rm DM}=1\,\mathrm{GeV\cdot cm^{-3}}$ ($2.63\times 10^{7}\,M_{\odot}\mathrm{\cdot kpc^{-3}}$).
@@
The numerical results of the photon–axion conversion probability in Fig. 1 demonstrate resonant conversion exactly at the locations predicted by Eq. (10), while the behavior reduces to the adiabatic limit away from resonance. This excellent agreement confirms that the Floquet analysis accurately describes the physics of the driven system. The conversion probability exhibits a distinct step-like growth, indicating that resonant conversion occurs within narrow spatial windows where the phase-matching condition is satisfied. The resonance positions depend solely on the jet’s plasma profile and the axion mass, independent of the DM density $\rho_{\rm DM}$. Thus, while a realistic non-uniform DM halo would modulate the amplitude of each ”step” according to local $\rho_{\rm DM}$ and $B$ values, the qualitative structure of multiple resonances remains unchanged.
@@
The spatial width of the parametric resonance $L_{\text{res}}$ can be estimated as $\left|4g_{n}/(\nabla_{r}\Delta_{n})\right|_{r_{\text{res}}}$. For the jet model considered here, the resonance width is approximately $2(nm_{a})^{-1}g_{a\gamma}B_{0}J_{n}(\alpha)r_{0}$. For the parameters used in Fig. 1, we obtain $L_{\text{res}}\sim 3\,\text{pc}$, in agreement with the numerical results. This finite width confirms that the resonance is robust against density fluctuations on scales smaller than $L_{\text{res}}$. Given that $L_{\text{res}}$ is two orders of magnitude smaller than the total jet length ($\sim\mathrm{kpc}$), plasma density variations within each resonance region are negligible. This separation of scales ensures that the parametric resonance remains robust even in highly inhomogeneous jet environments, where the global density gradient merely serves to sequentially satisfy discrete resonance conditions.
@@
The parametric resonance relies on the coherent oscillation of the axion field. The validity of this assumption depends on the coherence time of the DM field, $\tau_{\mathrm{coh}}\sim(m_{a}v^{2})^{-1}$, compared to the photon’s transit time through the resonance region, $t_{\mathrm{cross}}\sim L_{\mathrm{res}}$. For ultralight axions considered here with virial velocity $v\sim 10^{-3}$, the coherence time is $\tau_{\mathrm{coh}}\sim\mathcal{O}(10^{5})\,\text{years}$. In contrast, the spatial width $L_{\mathrm{res}}\sim 3\,\mathrm{pc}$ corresponds to a transit time $t_{\mathrm{cross}}\approx 10\,\text{years}$. The condition $t_{\mathrm{cross}}\ll\tau_{\mathrm{coh}}$ ensures that the DM background acts as a highly coherent driving force during the resonant interaction. Moreover, the fractional frequency spread $\delta m_{a}/m_{a}\sim 10^{-6}$ induces a negligible broadening of the resonance layer compared to its physical width determined by the jet geometry and coupling strength.
@@
Figure S1: ECDF of the photon-to-axion conversion probability $P_{\gamma\to a}$. The color coding indicates various turbulence correlation lengths $L_{fluc}$ for electron density (from $10^{-3}$ to $1\text{ pc}$), while solid and dashed lines distinguish between models with and without DM background, respectively. Each curve is derived from a sample size of $N=1000$. Note that the x-axis is presented in log-scale.
@@
To assess the robustness of the resonance mechanism, we introduce 50% stochastic electron density fluctuations across turbulence scales ranging from $L_{\rm fluc}=10^{-3}$ to $1$ pc. As shown in Fig. S1, the mean conversion probability in the DM background remains consistently 1-2 orders of magnitude higher than the standard baseline without DM, regardless of the correlation length. While small-scale fluctuations ($L_{\rm fluc}\ll L_{\rm res}$) are effectively self-averaged, larger-scale variations ($L_{\rm fluc}\sim 1$ pc) increase statistical variance by inducing local dephasing. Nevertheless, the empirical cumulative distribution function reveals that even the lower quantiles of these distributions significantly surpass the standard adiabatic case.
@@
Standard astrophysical emission mechanisms in blazars typically yield negligible circular polarization. However, parametric resonance breaks the degeneracy between left- and right-handed circular polarization states through the parity-violating coupling. This induces a non-zero net circular polarization, whose sign and magnitude are stochastic and depend on the phase $\varphi_{a}$ of the axion background at the resonance onset.
@@
We apply the formalism to constrain the axion–photon coupling $g_{a\gamma}$ using optical polarimetry of 3C 279. The DM density is modeled using a fuzzy DM profile, which consists of a solitonic core and an outer NFW envelope [Navarro:1995iw, Schive:2014dra] with the total mass $M_{h}\sim 10^{13}M_{\odot}$ [Ferrarese:2002ct]. The core density profile is adopted as $\rho_{c}[1+0.091(r/r_{c})^{2}]^{-8}$ with the core radius $r_{c}\simeq 1.6\,\mathrm{kpc}\left(m/10^{-22}\,\mathrm{eV}\right)^{-1}\left(M_{\rm h}/10^{9}\,M_{\odot}\right)^{-1/3}$ and density $\rho_{c}\simeq 2.9\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-3}(m/10^{-22}\,\mathrm{eV})^{2}(M_{\rm h}/10^{9}\,M_{\odot})^{4/3}$ [Schive:2014hza]. The transition radius between the two regions is taken to be $r_{t}\simeq 3r_{c}$. The concentration parameter of the NFW halo is determined using the concentration–mass relation at redshift $z=0.5$ [Dutton:2014xda]. This configuration yields negligible spatial gradients with $\mathcal{F}\lesssim 10^{-3}$ .
@@
Figure S2: KDE-estimated probability density of induced circular polarization for various $m_{a}$ ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$), obtained by uniformly sampling the initial axion phase. Shaded areas denote density profiles computed using Gaussian kernels with Scott’s rule bandwidths.

### v2

@@ Abstract
A direct phenomenological consequence of driven resonance crossings is the generation of polarization asymmetries between propagation modes. Repeated or stochastic crossings induce circular polarization with vanishing mean but nonzero variance. Because the axion-induced modulation enters with opposite signs for the two circular polarization states, the driven mixing mechanism is intrinsically polarization sensitive, even for initially unpolarized radiation. The variance of the circular polarization thus provides a direct observational signature of Floquet-driven photon–axion mixing.
Blazar 3C 279 as an illustration.– We validate this theoretical framework with full numerical solutions and demonstrate that the driven mixing mechanism can imprint characteristic signatures on photon polarization during propagation. As a concrete realization, we apply the formalism to optical polarimetry of the blazar 3C 279, with source-specific modeling details provided in the supplementary material.
To illustrate the potential observational impact of the mechanism, we adopt a statistical consistency criterion based on circular polarization fluctuations. Specifically, we require that the standard deviation of the axion-induced circular polarization, $\sigma_{\mathrm{std}}$, does not exceed the characteristic observational uncertainty of a given source. For 3C 279, we take $\sigma_{\Pi_{C}}\approx 0.30\%$ [26]. This choice reflects the fact that, while the axion field remains phase-coherent over observational timescales, its phase $\varphi_{a}$ is a priori unknown. As a result, the induced circular polarization does not correspond to a deterministic signal but instead manifests as a stochastic contribution with vanishing mean. In this situation, the variance provides a physically well-defined and observationally relevant measure of the effect.
Applying this criterion yields an illustrative bound on the axion–photon coupling, shown in Fig. 2 together with CAST constraints [5] for reference. For axion masses in the range $m_{a}\in[10^{-23},10^{-20}]\,\mathrm{eV}$, the resulting sensitivity reaches $g_{a\gamma}\lesssim 2\times 10^{-11}\,\mathrm{GeV}^{-1}$ for the adopted source parameters. While the numerical level of the bound depends on the specific noise properties of 3C 279, its qualitative structure is set by the underlying driven mixing dynamics.
In particular, the banded pattern in parameter space directly reflects the discrete Floquet resonance condition associated with the periodic axion background. As $m_{a}$ varies, the dominant contribution shifts between different resonance orders $n$, leading to alternating regions of enhanced and suppressed conversion. Near resonance, the driven mixing efficiently generates circular polarization, whereas between resonances the system approaches the adiabatic regime and the effect is correspondingly reduced. This characteristic structure is a direct consequence of the time-dependent Floquet driving and does not rely on fine-tuned phase alignment.
We emphasize that the bound presented here is an order-of-magnitude illustration, rather than a definitive competitive constraint. It demonstrates that the driven mixing mechanism can lead to observable, mechanism-specific polarization signatures at realistic sensitivity levels. More generally, the emergence of discrete resonance bands and stochastic circular polarization constitutes a robust phenomenological consequence of photon propagation in a coherently oscillating axion background.
@@
Figure 2: Illustrative bound on the axion–photon coupling $g_{a\gamma}$ derived from 3C 279 optical polarimetry. The shaded region indicates parameters excluded by the non-observation of excess circular polarization. The driven mixing mechanism gives rise to a distinctive band structure in parameter space, directly reflecting the discrete Floquet resonance condition $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$.
@@
Conclusions.– Wave propagation in a coherently oscillating background is intrinsically a periodically driven process. Photon propagation through axion dark matter therefore exhibits a driven resonant mode-conversion phenomenon that lies beyond the conventional static or adiabatic mixing paradigm. Resonant conversion occurs when the mismatch between photon and axion dispersion relations satisfies $\Delta_{\gamma}-\Delta_{a}\approx nm_{a}$, allowing coherent conversion even far from the standard level-crossing condition.
@@
This driven resonance represents a distinct class of propagation effects. Unlike conventional resonant conversion, which is localized to regions where plasma and axion masses coincide, it arises from temporal phase matching induced by the oscillating background and can persist over extended propagation distances. Small conversion amplitudes then accumulate coherently, leading to characteristic polarization signatures.
The effect depends on phase coherence along the propagation path. Rapidly varying environments suppress the resonance, while sufficiently smooth backgrounds allow driven conversion to operate effectively and, in some regimes, to dominate over static mixing contributions.
More generally, the framework developed here applies to wave propagation in coherently oscillating backgrounds beyond the specific axion model considered. Because the resonance relies on temporal periodicity rather than on the microscopic origin of the oscillations, analogous driven mode conversion may arise in multi-axion or other dark-sector scenarios.
Taken together, these results establish a systematic description of resonant wave propagation in time-dependent media and identify a previously overlooked driven regime of photon–axion mixing induced by coherent dark matter oscillations.
@@ Acknowledgements.
Acknowledgments.– The authors would like to thank the computing cluster of School of Fundamental Physics and Mathematical Sciences at Hangzhou Institute for Advanced Study, UCAS for providing the computational resources used in this work. This work is supported by the National Natural Science Foundation of China under Grants No. 12547110,12475065, 12250010, 12447131, 12447105, and 12575113.
@@
This Supplementary Material (SM) contains additional calculation and derivation in support of the results presented in this work. First, we give a full derivation of the Hamiltonian presented in the main text. Then, we present additional details regarding the derivation of the two-level systems. Next, we discuss the relation between our work and parametric instability. Finally, we present a numerical analysis and apply the formalism to the relativistic jet of blazar 3C 279.
@@
In this section, we provide the step-by-step derivation of the Hamiltonian presented in the main text.
@@
For a weakly magnetized cold plasma, the dielectric tensor is isotropic, $\bm{\epsilon}=(1-\frac{\omega_{\mathrm{pl}}^{2}}{\omega^{2}})\mathbb{I}$, where $\omega_{\mathrm{pl}}=\sqrt{4\pi\alpha n_{e}/m_{e}}$ is the plasma frequency and $\omega$ is the energy of the photon–axion system. Here, $\alpha$ is the fine-structure constant, $n_{e}$ and $m_{e}$ denote the electron number density and mass, respectively. In the radiation gauge, longitudinal plasma oscillations decouple from the propagating electromagnetic modes. In Fourier space, the induced plasma current therefore contributes an isotropic dispersion term $\mathbf{J}_{\mathrm{ind}}(\omega)=-\omega^{2}(\mathbb{I}-\bm{\epsilon})\mathbf{A}(\omega)=-\omega_{\mathrm{pl}}^{2}\mathbf{A}(\omega)$ to both photon polarizations, which can be absorbed into an effective photon mass.
@@
$\mathcal{M}(z)=\begin{pmatrix}\Delta_{\perp}&\Delta_{F}&0\\ \Delta_{F}^{*}&\Delta_{\parallel}&\Delta_{\mathrm{B}}\\ 0&\Delta_{\mathrm{B}}&\Delta_{a}\end{pmatrix}$. (S13)
Here, $\Delta_{\perp,\parallel}$ represent the photon dispersion for the two linear polarizations, accounting for both plasma effects and vacuum birefringence. In the systems considered here, vacuum QED effects are included in $\Delta_{\perp,\parallel}$ but remain subdominant to the plasma contribution, $\Delta_{\perp,\parallel}\simeq-\omega_{\mathrm{pl}}^{2}/(2\omega)$. The axion mass-induced dispersion is given by $\Delta_{a}=-m_{a}^{2}/(2\omega)$, while the photon–axion mixing is governed by $\Delta_{\rm B}=g_{a\gamma}B_{\mathrm{T}}/2$ with the transverse magnetic field $B_{\mathrm{T}}$. The $\Delta_{F}$ term originates from the axion-gradient-induced birefringence and is formally analogous to the Faraday rotation term. Explicitly, it is given by
@@
In the derivation above, we have neglected the absorption effects of photons, the contribution of the cosmic microwave background’s energy density to the dispersion relation, and the QED vacuum birefringence effect.
@@ II Effective Two-Level systems
where $\Delta_{\pm}=\Delta_{\gamma}\mp\Delta_{\text{DM}}\sin\phi$ represent the photon dispersion for the two circular polarizations. Neglecting the QED birefringence effect, the evolution for the photon state can be formally written as
@@
Repeating the Fourier harmonic expansion as above, one would also obtain the same resonance condition and Rabi oscillations. It has been noted that coherent axion background induces a periodic frequency modulation of photon propagation in Ref. [29]. However, in that work the modulation is treated as a phase modulation of a fixed propagation eigenmode, and physical observables are obtained after averaging over the axion oscillation. As a result, the associated sidebands are not promoted to independent dynamical degrees of freedom. This procedure effectively restricts the dynamics to a single Floquet sector and is equivalent to replacing the time-ordered evolution operator by one generated by an averaged Hamiltonian. Such treatments effectively average over the harmonic structure of the modulation and therefore do not capture the Floquet sidebands responsible for resonant mode mixing.
@@ III.2 Regime I: Parametric Instability (The Mathieu Limit)
where $\epsilon\sim g_{a\gamma}a_{0}m_{a}/\omega$. The defining feature of this regime is that the periodic axion background modulates the effective frequency squared of the oscillator. This modulation couples the positive-frequency ($e^{-i\omega t}$) and negative-frequency ($e^{+i\omega t}$) modes. When $m_{a}\simeq 2\omega$, the system enters an instability band where the field amplitude grows exponentially.  [^1] The instability of the photon–axion system—considering both axion DM and magnetic fields but omitting the plasma background—is explored in Ref. [28], where an instability condition $\sqrt{k_{\gamma}^{2}+m_{a}^{2}}+k_{\gamma}=nm_{a}$ is identified. This corresponds to the stimulated decay of the axion field into photon pairs, a genuine parametric instability where the background acts as an energy pump.
@@ IV Numerical analysis
To numerically verify the analytical framework, we solve the Eq. (S16) incorporating the mixing matrix $\mathcal{M}$ for a concrete astrophysical environment. We model the relativistic jet of blazar 3C 279, focusing on the unpolarized optical R-band photons ($E_{\gamma}\approx 1.9\,\mathrm{eV}$) with a representative parameter set: an axion mass $m_{a}=10^{-22}\,\mathrm{eV}$ and a coupling constant $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$. The radiation zone, located $r_{0}=0.5\,\mathrm{pc}$ from the central engine, is characterized by a magnetic field $B_{0}=0.5\,\mathrm{G}$ and electron density $n_{e,0}=5\times 10^{3}\,\mathrm{cm}^{-3}$ [1]. The jet is modeled as a conical outflow with a Doppler factor $\delta\simeq 13$ [1], assuming scaling profiles $B(r)\propto r^{-1}$ and $n_{e}(r)\propto r^{-2}$ over a total length $L_{\rm jet}=1\,\mathrm{kpc}$. To capture the fine-grained evolution of the mixing process, the calculation is performed over domain sizes ranging from $5\times 10^{-7}\,\mathrm{pc}$ to $10^{-3}\,\mathrm{pc}$.
@@
To isolate and illustrate the resonance structures dictated by the plasma profile, we adopt a uniform DM density $\rho_{\rm DM}=1\,\mathrm{GeV\cdot cm^{-3}}$ ($2.63\times 10^{7}\,M_{\odot}\mathrm{\cdot kpc^{-3}}$), and obtain the photon–axion conversion probability shown in Fig. 1. Fig. 1 demonstrates resonant conversion exactly at the locations predicted by Eq. (11), while the behavior reduces to the adiabatic limit away from resonance. This excellent agreement confirms that the Floquet analysis accurately describes the physics of the driven system. The conversion probability exhibits a distinct step-like growth, indicating that resonant conversion occurs within narrow spatial windows where the phase-matching condition is satisfied. The resonance positions depend solely on the jet’s plasma profile and the axion mass, independent of the DM density $\rho_{\rm DM}$. Thus, while a realistic non-uniform DM halo would modulate the amplitude of each ”step” according to local $\rho_{\rm DM}$ and $B$ values, the qualitative structure of multiple resonances remains unchanged.
@@
The spatial width of the Floquet resonance $L_{\text{res}}$ can be estimated as $\left|4g_{n}/(\nabla_{r}\delta_{n})\right|_{r_{\text{res}}}$. For the jet model considered here, the resonance width is approximately $2(nm_{a})^{-1}g_{a\gamma}B_{0}J_{n}(\alpha)r_{0}$. For the parameters used in Fig. 1, we obtain $L_{\text{res}}\sim 3\,\text{pc}$, in agreement with the numerical results. This finite width confirms that the resonance is robust against density fluctuations on scales smaller than $L_{\text{res}}$. Given that $L_{\text{res}}$ is two orders of magnitude smaller than the total jet length ($\sim\mathrm{kpc}$), plasma density variations within each resonance region are negligible. This separation of scales ensures that the resonance remains robust even in highly inhomogeneous jet environments, where the global density gradient merely serves to sequentially satisfy discrete resonance conditions.
@@
The Floquet resonance relies on the coherent oscillation of the axion field. The validity of this assumption depends on the coherence time of the DM field, $\tau_{\mathrm{coh}}\sim(m_{a}v^{2})^{-1}$, compared to the photon’s transit time through the resonance region, $t_{\mathrm{cross}}\sim L_{\mathrm{res}}$. For ultralight axions considered here with virial velocity $v\sim 10^{-3}$, the coherence time is $\tau_{\mathrm{coh}}\sim\mathcal{O}(10^{5})\,\text{years}$. In contrast, the spatial width $L_{\mathrm{res}}\sim 3\,\mathrm{pc}$ corresponds to a transit time $t_{\mathrm{cross}}\approx 10\,\text{years}$. The condition $t_{\mathrm{cross}}\ll\tau_{\mathrm{coh}}$ ensures that the DM background acts as a highly coherent driving force during the resonant interaction. Moreover, the fractional frequency spread $\delta m_{a}/m_{a}\sim 10^{-6}$ induces a negligible broadening of the resonance layer compared to its physical width determined by the jet geometry and coupling strength.
(a) Fluctuations in electron density
(b) Fluctuations in magnetic field
@@
Figure S1: Empirical cumulative distribution function (ECDF) of the photon-to-axion conversion probability $P_{\gamma\to a}$. The colors represent different fluctuation correlation lengths $L_{\rm fluc}$, ranging from $10^{-3}$ to 11 pc. Panels (a) and (b) correspond to stochastic fluctuations in the electron density and magnetic field, respectively. Solid and dashed lines distinguish between driven Floquet mixing with oscillating DM background and adiabatic conversion with static DM background. Each curve is generated from $N=1000$ random realizations. Note that the $x$-axis is on a logarithmic scale.
The phenomenological impact of this mechanism depends on the degree to which phase coherence is maintained along the propagation path. Strong gradients or rapid environmental variations can disrupt the driven mixing and suppress the resonance, while sufficiently smooth backgrounds allow the periodically driven dynamics to operate effectively. In such regimes, the driven contribution to photon–axion mixing can be comparable to, or exceed, that expected from static conversion alone.
@@
To assess the robustness of the resonance mechanism, we investigate the effects of stochastic fluctuations in the electron density and the magnetic field separately. The propagation path is partitioned into discrete domains of length $L_{\rm fluc}$. For each scenario, we introduce a random bias by scaling the baseline model value by a coefficient $\xi$, uniformly sampled from $[0.5,1.5]$, thereby representing $50\%$ amplitude fluctuations. As shown in Fig. S1, even with these 50% amplitude variations, the mean conversion probability under Floquet-driven mixing remains consistently 1–2 orders of magnitude higher than in the standard static scenario across all considered correlation lengths. While small-scale fluctuations ($L_{\rm fluc}\ll L_{\rm res}$) are effectively self-averaged, larger-scale variations ($L_{\rm fluc}\sim 1$ pc) increase statistical variance by inducing local dephasing. Nevertheless, the empirical cumulative distribution function (ECDF) reveals that even the lower quantiles of these distributions significantly surpass the standard adiabatic case.
@@
Standard astrophysical emission mechanisms in blazars typically yield negligible circular polarization. However, the Floquet resonance breaks the degeneracy between left- and right-handed circular polarization states through the parity-violating coupling. This induces a non-zero net circular polarization, whose sign and magnitude are stochastic and depend on the phase $\varphi_{a}$ of the axion background at the resonance onset.
This stochastic circular polarization arises from coherent, unitary mode mixing and is therefore qualitatively distinct from polarization generated by absorption or scattering processes. The magnitude of the variance is directly controlled by the resonance strength, thereby providing a direct probe of the driven mixing mechanism. Importantly, this polarization signature does not rely on fine-tuned source properties or detailed environmental modeling. It follows directly from the existence of Floquet resonances in photon propagation through a coherent axion background.
@@
We apply the formalism to constrain the axion–photon coupling $g_{a\gamma}$ using optical polarimetry of 3C 279. The DM density is modeled using a fuzzy DM profile, which consists of a solitonic core and an outer NFW envelope [30, 42] with the total mass $M_{h}\sim 10^{13}M_{\odot}$ [20]. The core density profile is adopted as $\rho_{c}[1+0.091(r/r_{c})^{2}]^{-8}$ with the core radius $r_{c}\simeq 1.6\,\mathrm{kpc}\left(m/10^{-22}\,\mathrm{eV}\right)^{-1}\left(M_{\rm h}/10^{9}\,M_{\odot}\right)^{-1/3}$ and density $\rho_{c}\simeq 2.9\times 10^{6}\,M_{\odot}\,\mathrm{kpc}^{-3}(m/10^{-22}\,\mathrm{eV})^{2}(M_{\rm h}/10^{9}\,M_{\odot})^{4/3}$ [43]. The transition radius between the two regions is taken to be $r_{t}\simeq 3r_{c}$. The concentration parameter of the NFW halo is determined using the concentration–mass relation at redshift $z=0.5$ [17]. This configuration yields negligible spatial gradients with $\mathcal{F}\lesssim 10^{-3}$ .
@@
Figure S2: Probability density of induced circular polarization for various axion masses $m_{a}$ (assuming $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV}^{-1}$), estimated via Kernel Density Estimation (KDE) by uniformly sampling the axion phase $\varphi_{a}\in[0,2\pi]$. The shaded profiles are generated using Gaussian kernels with Scott’s rule bandwidths. The transition from single-peaked to double-peaked distributions illustrates the varying sensitivity of the conversion process to the axion phase across different mass regimes.